The density of ethanol is 0.789g/ml. How many qts would 0.55 kg occupy?
Answer #1

Volume of ethanol = Mass of ethanol / Density of ethanol = 0.55 * 1000 / 0.789 = 697.085 mL

1 qt = 946.353 ml

Volume of ethanol in qt = 697.085 / 946.353 = 0.737 qt